# Exploring human-SAV interaction using large language models: The impact of psychological ownership and anthropomorphism on user experience

### v1

@@ 3.2 SAV prompt design
    Standard + [Psychological Ownership] prompt
    The second group incorporated strategies to foster a sense of psychological ownership by following the three PO routes introduced in Section 2.1 – exercise of control, intimate knowledge, and investment of the self. The large language model was instructed to ensure users perceive the SAV as “their own private autonomous vehicle” by emphasizing control over vehicle operations, to share intimate knowledge about the SAV, and encourage users to invest themselves in its use. The goal was to enhance user attachment and engagement, and develop a strong sense of psychological ownership while users interacted with the vehicle.
@@
    Standard + [Anthropomorphism] prompt
  4. 4.
    Standard + [Psychological Ownership] prompt + [Anthropomorphism] prompt
Each prompt was designed to reflect the unique objectives of the study and was implemented as part of the SAV’s conversational UI. The prompts were iteratively trialed and refined to ensure they conveyed the SAV’s intended communication style, functionality, and user experience. The detailed descriptions and wording of the prompts can be found in the Appendix A.1. The inclusion of these variations allowed for a comparative analysis of how different prompts influence user engagement, satisfaction, and psychological ownership.
3.3 User study design
A user study was conducted with ethical approval from the Monash University Human Research Ethics Committee (MUHREC) [^2] Project ID: 40485. Participants were invited to interact with all four SAV groups (in random order). A list of fifteen sample user inputs (e.g., “Adjust the seat, I want to lie down”) was provided, but participants were also informed that they were free to interact with the SAVs in any way they desired, including making requests or asking questions. After each interaction, participants were asked to fill out a survey about their interaction experience covering the following aspects: their feelings of psychological ownership, the perceived anthropomorphism of the SAV, quality of service, their disclosure tendency, polarity (i.e., whether the responses of the SAVs were positive or negative), subjectivity (i.e., whether the responses of the SAVs were subjective or objective), enjoyment, and SAV acceptance. The questionnaire items used are summarized in Table 1. Qualitative data was collected from open-ended questions about additional features or experiences with the vehicle that would contribute to psychological ownership, the benefits of sharing personal information, and interview responses about participants’ perceived psychological ownership. The user studies were held from April 30, 2024, to July 2, 2024. In total, 50 participants were invited, and 2,136 chats (a question-answer interaction) were collected. In this research project, the targeted population was selected to be young and highly educated, as this group may represent potential early adopters of technologies like SAVs due to their openness to innovation and higher levels of education (Guo et al., 2025).
Table 1: Questionnaire Items and Sources.
Questionnaire items Modified from Psychological Ownership (PO)—Measured from 1 (Strongly Disagree) to 7 (Strongly Agree) PO1: I feel like the vehicle is mine Lee et al. (2019); Li et al. (2023) PO2: I feel a very high degree of personal ownership of the vehicle PO3: Interacting with the SAV interface enhances my feeling of ownership over the vehicle PO4: I feel like I own the vehicle Anthropomorphism (A)—Measured from 1 (Strongly Disagree) to 7 (Strongly Agree) A1: I believe that the SAV has its own decision-making power Pelau et al. (2021); Chi et al. (2021) A2: I believe that the SAV has its own personality A3: I believe that the SAV has its own preferences and mood A4: I believe that the SAV has its own emotions A5: I believe that the SAV device can feel compassion Quality of Service (QoS)—Measured from 1 (Strongly Disagree) to 7 (Strongly Agree) QoS1: I am satisfied with the SAV service. Zhang and Patrick Rau (2023) QoS2: I find communication with the SAV to be pleasant. QoS3: I would recommend the SAV to others. Disclosure Tendency (DT)—Measured from 1 (Strongly Disagree) to 7 (Strongly Agree) DT1: I am willing to disclose my daily route with the SAV. Zhang and Patrick Rau (2023) DT2: I am willing to disclose my tastes in music with the SAV. DT3: I am willing to disclose my feelings, including negative emotions like anxiety, with the SAV. Polarity (Pol)—Measured from -1 (Negative) to 1 (Positive) Pol: Overall, how do you rate the SAV responses, negative or positive? By definition, Wang et al. (2014) Subjectivity (Sub)—Measured from 0 (Objective) to 1 (Subjective) Sub: Overall, do you think the SAV responses are objective or subjective? By definition, Wang et al. (2014) Perceived Enjoyment (PE)—Measured from 1 (Strongly Disagree) to 7 (Strongly Agree) PE1: How fun do you feel your interaction with this SAV was? Pelau et al. (2021) PE2: Do you think “fun” is a factor that would influence your intention to use SAVs? Behavioral Intention to Use (BI)—Measured from 1 (Strongly Disagree) to 7 (Strongly Agree) BI1: Overall, would you use an SAV if given the chance? Lee et al. (2019); Dai et al. (2021) BI2: Would you likely use SAVs repeatedly in the future based on your experience today? BI3: Do you want to interact with this/your SAV again? *Note: item PE1 is measured from 1 (Not fun at all) to 7 (Extremely fun)
3.4 Hypotheses and statistical analysis
3.4.1 Descriptive statistics
Means and standard deviations (Tables 3) were calculated for all constructs within the survey to examine differences in participants’ perceptions of the selected psychological factors and their user experiences. A pie chart (Fig. 4) was plotted to visualize the preferred SAV choices.
3.4.2 Favorite SAVs
To investigate the effects of prompts on participants’ user experiences with the four types of SAVs, we first tested whether participants’ favorite SAVs varied significantly. If all SAVs were equally preferred, it could indicate that the tailored prompts utilized to develop the four SAV groups did not achieve the anticipated effect on users’ experiences or choices. This resulted in the formulation of our initial hypothesis:
  • H1: The favorite SAVs are different across all participants.
Although the order of SAV presentations was randomized to minimize any ordering effect on participants’ perceptions, it is essential to examine whether the presentation order had any influence on their preferences. For instance, as participants engage with the SAV UIs over time, they may develop a heightened sense of psychological ownership due to increased familiarity, which could impact the efficacy of the prompts. In light of the aforementioned considerations, we put forth a second hypothesis:
  • H2: There is a relationship between the SAV order and the favorite SAV.
The user study employed three distinct anthropomorphic characteristics for the SAVs, which were as follows: “cool and sophisticated”, “friendly and engaging”, and “sassy and tired”. Participants selected their preferred anthropomorphic style before interacting with the SAVs and indicated their favorite SAV at the end of all interactions. It is important to ascertain whether participants’ preferences for particular SAVs were shaped by the tailored prompts or by the SAVs’ anthropomorphic attributes. In other words, participants may exhibit a systematic preference for one SAV over others based on their perception of its interaction style. This leads to our third hypothesis, which is as follows:
  • H3: There is an association between the selected anthropomorphism (e.g., friendly or cool) and the favorite SAV.
To test these hypotheses, non-parametric tests were used. Specifically, the Chi-Square Test and Fisher’s Exact Test were applied depending on whether the expected frequency for each category was greater or less than 5, respectively.
3.4.3 Prompt effect
The effects of the designed prompts were examined by comparing the survey responses collected after each interaction with the SAVs. Cronbach’s alpha values were first assessed to verify the internal consistency of the questionnaire items. Next, the Friedman test, a non-parametric alternative to repeated measures ANOVA that is robust to departures from normality (Ramachandran and Tsokos, 2015), was conducted to determine whether there were significant differences between the four SAVs. Specifically, the test was applied to psychological ownership (PO), anthropomorphism (A), quality of service (QoS), disclosure tendency (DT), polarity (Pol), subjectivity (Sub), enjoyment of interaction (PE), and behavioral intention to use (BI). Any significant variation in these constructs would indicate that the prompts influenced participants’ perceptions and experiences, consistent with the following overarching hypothesis:
  • H4: A significant difference exists among the four SAV groups across the measured variables.
To investigate how PO prompts and A prompts influence user perceptions and experiences, a series of one-tailed Wilcoxon Signed-Rank tests were employed at the item level. In addition, Holm–Bonferroni corrections were applied within each family of pairwise comparisons to control the family-wise error (Armstrong, 2014).
Psychological ownership prompt effect
It was first hypothesized that integrating PO prompts in an SAV interface would enhance users’ sense of ownership towards the SAV. Specifically, it was hypothesized that:
  • H5.1: SAV2 (integrated with PO prompts) > SAV1 (Baseline) on PO items (PO1–PO4).
  • H5.2: SAV4 (both PO and A prompts) > SAV3 (integrated with A prompts) on PO items (PO1–PO4).
Building on Section 2.1, it was further proposed that PO prompts might affect broader user experience factors, which includes QoS, DT, sentiment of the SAV responses (Polarity and Subjectivity), PE, and BI:
  • H6.1: SAV2 (PO prompts) > SAV1 (Baseline) on these user experience metrics.
  • H6.2: SAV4 (both PO and A prompts) > SAV3 (A prompts) on these user experience metrics.
Anthropomorphism prompt effect
It was expected that incorporating anthropomorphic prompts in the SAV interface could enhance users’ perception of the vehicle’s anthropomorphic qualities. It was hypothesized that:
  • H7.1: SAV3 (A prompts) > SAV1 (Baseline) on A items (A1–A5).
  • H7.2: SAV4 (both PO and A prompts) > SAV2 (PO prompts) on A items (A1–A5).
Similarly, building on Section 2.2, it was hypothesized that A prompts would also enhance these user experience dimensions:
  • H8.1: SAV3 (A prompts) > SAV1 (Baseline) on user experience metrics.
  • H8.2: SAV4 (both PO and A prompts) > SAV2 (PO-only prompts) on user experience metrics.
Combined prompt effect
Finally, an exploratory analysis examined the combined effects of PO and A prompts by comparing SAV4 (both PO and A prompts) against SAV1 (Baseline). It was expected to see that SAV4 outperforms SAV1 in all PO, A, and user experience metrics aspects.
  • H9.1: SAV4 (both PO and A prompts) > SAV1 (Baseline) on PO perception.
  • H9.2: SAV4 (both PO and A prompts) > SAV1 (Baseline) on A items (A1–A5).
  • H9.3: SAV4 (both PO and A prompts) > SAV1 (Baseline) on user experience metrics.
3.4.4 Qualitative analysis
From the collected open-ended responses and interview data, seven themes were identified. Under each theme, various codes were developed to represent the aspects that participants mentioned as influencing the development of their psychological ownership when interacting with the SAVs. The themes and codes are as follows:
  • Communication Style: Short and concise, Long and detailed, Interactive and engaging, Verbose and repetitive, Guidance and recommendations, Use of “I”, Use of “we”;
  • Communication Tone: Professional, polite, and formal, Kind and warm, Neutral;
  • Anthropomorphism: Human-like, AI-/GPT-like, Robot-like, Machine-like;
  • Performance: Meeting user’s needs, Efficiency, Care and support, Accuracy, Intelligence and conversation memory, Perceived ease of use, Taking initial actions;
  • Psychological Ownership Routes: Personalization, Familiarity, intimacy and privacy, Feeling of control, Time and number of interactions;
  • Legal Ownership: Fact of shared car, Like a taxi driver/service;
  • Sentiment: Objective.
The effects of each code on participants’ perceptions of psychological ownership were analyzed, noting whether each influence was positive or negative, and the frequency of occurrences was recorded. The descriptions and sample responses illustrating each code are provided in Table LABEL:tab:prompt_code_summary.
@@ 3.4.4 Qualitative analysis
Table 2: Summary of Communication Styles and Effects in SAV Interactions. The corresponding SAV ID that was mentioned in the response examples is indicated in brackets. If no SAV ID is specified, it indicates a general comment that applies to all four SAVs.

### v2

@@ VI Conclusion and future research directions
    Providing users with intimate knowledge about this SAV;
@@
    Encouraging users to invest themselves in this SAV.
S1-C Prompt G3-C – Standard + [Anthropomorphism] Cool and sophisticated
  • Media: Listen to media and adjust your playback preferences;
The following phase is replaced for Friendly and engaging SAV
The following phase is replaced for Sassy and tired SAV
S1-D Prompt 4C – Standard + [Psychological Ownership] + [Anthropomorphism] Cool and sophisticated
  • Media: Listen to media and adjust your playback preferences;
TABLE S12: Primary user request type labels and descriptions assigned to each conversational exchange in the SAV dataset. Only the primary task was labeled per exchange, considering both the immediate request and the conversational context.
Interaction Type Label Description
Navigation & Motion Control Plan, modify, or query the route and ETA, and issue real-time driving commands (speed changes, lane shifts, stop-overs, overtakes).
Cabin Environment Adjust any in-cabin setting, such as temperature, lighting, seat position, windows, scents to meet the passenger’s comfort needs.
Entertainment Control Control music, radio, podcasts, video, and other infotainment content.
System Capabilities & Policies Explain the vehicle’s capabilities, status, maintenance history, pricing rules, or legal/operational constraints.
External Communication Place calls, send messages or emails, make bookings, and interact with external contacts or services.
Well-being Support Provide empathy, emotional support, or motivational dialogue to assist the passenger’s well-being and emotional state.
Social Chat Engage in casual conversation, greetings, jokes, or other non-task-oriented small talk.
Safety & Security Perform or request actions such as locking/unlocking, high-beam control, emergency stop to relieve perceived risks.
Alerts & Reminders Create, schedule, or deliver timers, wake-ups, arrival alerts, and other proactive notifications.
Personalization & Profiles Save, load, or modify user profiles and preferred settings.
Local Context Provide real-time location, time, temperature, weather, or recommendations on nearby points of interest and events.
  1. 1.
  2. 2.
    Providing users with intimate knowledge about this SAV;
  3. 3.
    Encouraging users to invest themselves in this SAV.
Appendix S2 Human-SAV Interaction type labels
Table S12 summarizes the user request type labels and their corresponding descriptions. Label assignment was based primarily on the user’s request, but also took into account the overall context of each conversational exchange pair, as well as preceding interactions when relevant (e.g., for follow-up requests). Because the aim was to provide a concise overview of user request types, only the primary task was labeled for exchanges that contained multiple requests (i.e., each exchange received a single label). For example, in the request, “i am unhappy today, could you pls tell me some jokes?”, both Well-being Support and Social Chat are applicable; however, since the user’s main concern was their emotional state and the jokes were requested as a means of support, the exchange was labeled as Well-being Support.
Note that the human–SAV conversational interaction data analyzed here are drawn from an open-source dataset, as detailed in Section III-D Data availability. Researchers interested in more granular analyses are encouraged to extend the taxonomy, for instance, by assigning multi-level task labels to individual chats where appropriate.
Appendix S3 Mean and standard deviation results for factor items by participants’ favorite SAV
Table S4 provides the standard deviation values corresponding to the mean survey results presented in Table S13, which offer additional insights into the variability of participants’ responses.
Appendix S4 Wilcoxon Signed-Rank test results
The one-tailed Wilcoxon Signed-Rank test results are summarized in Tables S15-S18. Holm–Bonferroni corrections were applied within each family of pairwise comparisons. The significance check was based on the p-adjusted values.
TABLE S13: Mean survey results based on the selected favorite SAV. The highest mean value for each survey item across the four groups is highlighted in bold.
((a))
((b))
((c))
((d))
TABLE S14: Standard deviation of factor items based on the selected favorite SAV. The highest standard deviation value for each survey item across the four groups is highlighted in bold.
((e))
((f))
((g))
((h))
TABLE S15: Wilcoxon Signed-Rank Test results examining the effect of psychological ownership prompts on user experience.
Item Comparison Statistic p-value p-adjusted Significant
DT3 SAV2 >> SAV1 207.0 0.5965 0.9602 $\times$
Pol SAV2 >> SAV1 363.0 0.3201 0.9602 $\times$
Sub SAV2 >> SAV1 344.0 0.4342 0.9602 $\times$
BI2 SAV2 >> SAV1 256.5 0.1973 0.7894 $\times$
DT3 SAV4 >> SAV3 286.5 0.1268 0.6338 $\times$
Pol SAV4 >> SAV3 591.5 0.0173 0.1211 $\times$
Sub SAV4 >> SAV3 469.5 0.0057 0.0453
BI2 SAV4 >> SAV3 360.0 0.0753 0.4518 $\times$
TABLE S16: Wilcoxon Signed-Rank Test results for examining the effect of anthropomorphism prompts on user experience.
Item Comparison Statistic p-value p-adjusted Significant
DT3 SAV3 >> SAV1 282.0 0.1504 0.6017 $\times$
Pol SAV3 >> SAV1 371.0 0.1791 0.6017 $\times$
Sub SAV3 >> SAV1 349.5 0.2875 0.6017 $\times$
BI2 SAV3 >> SAV1 262.5 0.2673 0.6017 $\times$
DT3 SAV4 >> SAV2 266.5 0.0300 0.1799 $\times$
Pol SAV4 >> SAV2 544.0 0.0053 0.0371
Sub SAV4 >> SAV2 442.0 0.0019 0.0156
BI2 SAV4 >> SAV2 292.0 0.0502 0.2511 $\times$
TABLE S17: Wilcoxon Signed-Rank Test results for examining the effect of anthropomorphism prompts on user perception.
Item Comparison Statistic p-value p-adjusted Significant
A1 SAV3 >> SAV1 417.0 0.0190 0.0570 $\times$
A2 SAV3 >> SAV1 394.0 0.0017 0.0137
A3 SAV3 >> SAV1 320.0 0.0344 0.0688 $\times$
A4 SAV3 >> SAV1 313.5 0.0046 0.0265
A5 SAV3 >> SAV1 360.5 0.2254 0.2254 $\times$
A1 SAV4 >> SAV2 419.0 0.0015 0.0132
A2 SAV4 >> SAV2 342.5 0.0031 0.0218
A3 SAV4 >> SAV2 309.0 0.0071 0.0284
A4 SAV4 >> SAV2 277.5 0.0006 0.0063
A5 SAV4 >> SAV2 276.5 0.0044 0.0265
TABLE S18: Exploratory analysis results for combined prompt effects, comparing SAV4 with SAV1.
Item Statistic p-value p-adjusted Significant
A1 444.5 0.0015 0.0115
A2 383.0 0.0002 0.0014
A3 361.0 0.0039 0.0155
A4 392.5 0.0021 0.0123
A5 369.0 0.0242 0.0673 $\times$
DT3 357.0 0.0377 0.0673 $\times$
Pol 566.0 0.0020 0.0123
Sub 603.0 0.0014 0.0115
BI2 328.5 0.0224 0.0673 $\times$
Appendix S5 Full Coding Scheme for Communication Styles in SAV Interactions
@@
TABLE S19: Summary of Communication Styles and Effects in SAV Interactions. The corresponding SAV ID that was mentioned in the response examples is indicated in brackets. If no SAV ID is specified, it indicates a general comment that applies to all four SAVs.